What two integers have a product of 64 and a sum of -20?
Answer by checkley75(3666)

X+Y=-20 X=-20-Y &
X*Y=64.
NOW SUBSTITUTE (-20-Y) FOR X IN THE SECOND EQUATION & SOLVE FOR Y
(-20-Y)*Y=64
-20Y-Y^2=64
Y^2+20Y+64=0
(Y+16)(Y+4)=0
Y=16=0
Y=-16 ANSWER. THEN X=-4
Y+4=0
Y=-4 ANSWER. THEN X=-16
PROOF
-16-4=-20
-20=-20 &
-4*-16=64
64=64